Shift summary, Owl 23 Dec 99

Yee Bob Hsiung (hsiung@fnal.gov)
Thu, 23 Dec 1999 07:13:47 -0600

Shift Summary for Owl shift, 23 Dec 99

On Shift: Arun T, BobH

Hours of beam available at all intensities: 6.6
Hours of beam available at nominal intensity: 6.6
Hours of E799 data (special runs not included): 6.2
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: L3 fraction alarm, stop/start runs,
changing tapes and a few kfrends.

========================================================================
Runs ended this shift:
Run 15268 - 17.4M events - BEAM_799

Tapes brought to FCC this shift: UP*38-43, UPS080-081
UPH044, UPK032, UPF061

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E799 datataking

CsI gains continue changing while the blockhouse being heated by
blockhouse heater. (CsI-top = 18.6 C at the end of shift)

Trigger latch crate 8 is bypassed which causes dcor eff monitoring
goes to pot and many empty trigger histograms and non-working hypmon
plots etc.
One hot kq latch bit in DC3 for run 15269.

One L3 accept fraction alarm lead us to stop run 15269, but hung, then
restart new run. Sydney's voice won't shut up until end oof run 15270.

Otherwise, smooth running.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15269 01:15 Mount onsplit tape UPF062.
01:50 " " UPE017.
01:50 ktevana errors, kfrend.
kq hot bit in DC3 which affects ytf too.
01:56 L3 accept fraction alarm, stop run, but it hung.
02:08 updated constants DCT0268
15270 1 02:25 start run
02:58 Mount onsplit tape UPX017
04:30 tape full, run stopped.
15272 1 04:53 start run with new tapes.

========================================================================
Follow up Notes on Previous Shift
- Installed new T0's.
- Watched channel 968. It has a 6 count ped shift now.

========================================================================
Items to Follow up on Next Shift

- Take as much good data as possible.

*** END of REPORT ***